An airplane can fly 660 miles with the wind in the same amount of time it takes to fly 540 miles against the same wind. If the air speed of the plane in still air is 150 miles per hour, what is the speed of the wind?
Answer by mananth(16949)

plane speed = 150 kmph
wind speed = x kmph
speed against wind= 150-x
speed with wind 150+x
Distance against wind 540 km
Distance with tail wind 660 km
t=d/r time upstream = time downstream
540/(150-x)=660/(150+x)
540(150+x)=660(150-x)
81000+540x=99000-660 x
660x+540x=18000
1200x=18000
/1200
x=15mph wind speed